OpenStudy (kva1992):

sin(2A+30°) = -1/2 how do you find the solution for that?

OpenStudy (kva1992):

at first i thought i could use the sum difference identity but that does not seem like it would work because i do not know the value of 2A

OpenStudy (xapproachesinfinity):



OpenStudy (xapproachesinfinity):

look when does sin get negative

OpenStudy (xapproachesinfinity):

No need for sum difference identity

OpenStudy (xapproachesinfinity):

consider 2A+30=X solve for x then replace that with 2A+30m how is that

OpenStudy (kva1992):

it goes neg in q3 and q4

OpenStudy (xapproachesinfinity):

ok then fo what values does sin equal 1/2

OpenStudy (kva1992):

so would i just get what sin of 1/2 equals as theta hat ?

OpenStudy (kva1992):

and multiply that by whatever to get to either the 3 or 4th quad?

OpenStudy (xapproachesinfinity):

yes!

OpenStudy (kva1992):

so 7pi/6 or 11pi/6?

OpenStudy (xapproachesinfinity):

good! that's for big X what about 2A+30

OpenStudy (kva1992):

so it would be 210° + 360k where k represents all real numbers?

OpenStudy (xapproachesinfinity):

I think now everything is solved already ^^

OpenStudy (kva1992):

and it would be 330° +360k where k represents all real numbers?

OpenStudy (xapproachesinfinity):

do you need the answer in degree or radians?

OpenStudy (kva1992):

he said if it starts with degrees leave it as degrees

OpenStudy (xapproachesinfinity):

ok then leave, don't you need to solve for A?

OpenStudy (xapproachesinfinity):

because that's the answer for 2A+30

OpenStudy (kva1992):

btw he said the answers are 90° + 180°k and 150° 180°k where k is all real numbers

OpenStudy (xapproachesinfinity):

solve for A and see what happens

OpenStudy (xapproachesinfinity):

you should get those two answer I guess

OpenStudy (kva1992):

hmmm lemme check

OpenStudy (xapproachesinfinity):

ok

OpenStudy (kva1992):

nope neither gives me those answers

OpenStudy (xapproachesinfinity):

hmm let me see

OpenStudy (kva1992):

okie i pretty sure im not doing anything wrong but i am human after sorta our nature to get stuff wrong lol cant learn without making mistakes :)

OpenStudy (xapproachesinfinity):

Check again, I got those answers

OpenStudy (xapproachesinfinity):

did you solve 2A+30=210+ 360k

OpenStudy (kva1992):

hold up i just realized something lol 2A = 90 because a =45 degrees right?

OpenStudy (kva1992):

nope i dont think i set them equal to each other. mind if i ask why we are setting them equal to each other?

OpenStudy (xapproachesinfinity):

we started with sin(2A+30)=-1/2 then we said that 2A+30 = 210 + 360k or 2A+30= 330 +360k right?

OpenStudy (xapproachesinfinity):

we need to solve for angle A

OpenStudy (kva1992):

oh ya alright im with you so far

OpenStudy (kva1992):

oh so in order to do that you have to alright i think i get it

OpenStudy (xapproachesinfinity):

we just solved for 2A+30 we need to solve for A

OpenStudy (xapproachesinfinity):

Just think of 2A+30 as X in the beginning then after replace X with 2A+3) and solve for A

OpenStudy (xapproachesinfinity):

does that make sense?

OpenStudy (kva1992):

i think it does

OpenStudy (kva1992):

lemme double check though

OpenStudy (xapproachesinfinity):

Ok good

OpenStudy (kva1992):

the reason we got 210 and 330 is because sin = -1/2 in Q3 and Q4 right

OpenStudy (xapproachesinfinity):

yup, right!

OpenStudy (kva1992):

and once we find that we just set the equation they gave us equal to Q3 and Q4 and solve correct?